Show that in the definition of a vector space v the condition about existence of additive inverse can be replaced with the condition:0v=v for all v€V


1
Expert's answer
2021-05-07T11:35:36-0400

Given that V(F) is a vector space


"0.v=(0+0).v=0.v+0.v"


"\Rightarrow 0.v+0=0.v+0.v" Where 0 is the identity element in v


"\Rightarrow 0.v=0" (by left cancellation law)


hence "0.v=0 \text{ for all } v\in V"